40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 5 09:20:44 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_Chamber_optical_layout.pdf
    Reply  Thu Aug 5 12:39:56 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_in-air_setup.pdf
       Reply  Thu Aug 5 14:29:42 2021, Stephen, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber IMG_9685.JPGIMG_9687.JPGrouting_markup_of_IMG_9688.pngIMG_9682.JPG
          Reply  Tue Aug 24 15:01:25 2021, Stephen, Laser, 2um Photodiodes, 2um PD in new cryo chamber 
             Reply  Thu Aug 26 14:31:46 2021, Aidan, Stephen, Laser, 2um Photodiodes, Incident power calibration of 2um PD in new cryo chamber IMG_4420.jpgREF_PD_POW_CAL_cryo.pdf
                Reply  Fri Aug 27 17:57:30 2021, Stephen, DailyProgress, 2um Photodiodes, Start of PD pumpdown  
                   Reply  Sat Aug 28 07:01:21 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  20210828_-_IMG_4445.jpg20210828_-_IMG_4443.jpg
                      Reply  Sun Aug 29 11:17:43 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  Screen_Shot_2021-08-29_at_11.16.55_AM.pngScreen_Shot_2021-08-29_at_11.22.03_AM.png
                         Reply  Mon Aug 30 08:04:52 2021, Aidan, DailyProgress, 2um Photodiodes, Monday morning observation of JPL A1 PD cool down  Screen_Shot_2021-08-30_at_8.04.15_AM.png
Message ID: 2656     Entry time: Sun Aug 29 11:17:43 2021     In reply to: 2655     Reply to this: 2657
Author: Aidan 
Type: DailyProgress 
Category: 2um Photodiodes 
Subject: Start of PD pumpdown  

Sunday morning update:

11:17AM - Temperature of photodiode is ~86K

Also shown is the photodiode current time series. The laser is being ramped from minimum to maximum every 50s. The span of the signal increases as temperature is decreased. The response is almost nil at room temperature. Then it peaks around 140K and is clearly reduced by the time we get below 100K. (Caveat: the beam was initially aligned onto the PD at room temperature but there is currently no auto-alignment. The diode is quite large, 1.5mm sq, but we can;t rule out that the beam is moving off the PD as temperature decreases - so we can only put an lower limit on QE vs temperature from this data).

Attachment 1: Screen_Shot_2021-08-29_at_11.16.55_AM.png  366 kB  Uploaded Sun Aug 29 12:18:02 2021  | Hide | Hide all
Screen_Shot_2021-08-29_at_11.16.55_AM.png
Attachment 2: Screen_Shot_2021-08-29_at_11.22.03_AM.png  293 kB  Uploaded Sun Aug 29 12:23:59 2021  | Hide | Hide all
Screen_Shot_2021-08-29_at_11.22.03_AM.png
ELOG V3.1.3-